Sunday, July 8, 2012

Day 42, 7/8/2012 -- A wretched hive of scum and villainy

The EBEX bolo system power crates, that is -- designed years ago, they are probably the worst-laid-out piece of major equipment on the experiment. One of them is, anyway -- power crate 1, which houses DC/DC converters for half of the bolo system, the housekeeping system, the bolo side network, and a general housekeeping board which, among other things, controls various parts of the HWP system). Needless to say, this crate is very crowded, which means that we try as hard as we possibly can to not mess with it, lest we be forced to delve into the bowels of the crate.

Well, we were forced to. In order to try and solve a nagging DfMUX board power issue, we had to swap out the DC/DC voltage setting resistors for the +/- 5.8V power system, which entailed opening up the crate (as much as we could) and swapping out these resistors to drop the voltage by about 200mV. Along the way, we found that a power return wire for one of the DC/DCs was no longer connected and was subsequently fixed.

Power crate fixes, though, happened at the end of the day (we are just wrapping up now, at midnight). During the day, we were able to take spectral measurements for two 410GHz detectors. Kate and Kevin are now on the 'early' shift (7AM - ~4PM), so they could get some time with the DfMUX system without me and Kyle trying to get data.

The gondola crew spent much of today doing gyro tests, calibrating each individual gyro (there are six total, two full three-axis redundant gyro boxes) by swapping each gyro to a position where it would sense elevation motion and doing careful elevation slews. Britt worked on a little electronics board for the solar power system, and Michele worked on straightening out a small ding in our spare actuator's outer housing tube.


No comments:

Post a Comment

Note: Only a member of this blog may post a comment.